OpenStudy (anonymous):

The volume of a cube is increasing at a rate of 10cm^3/min. How fast is the surface area increasing when the length of the edge is 30cm.

OpenStudy (aravindg):

whats the formula for volume of cube?

OpenStudy (anonymous):

V=a^3

OpenStudy (aravindg):

here rate is given so what u do ?

OpenStudy (anonymous):

find the derivative with respect to the rate? ;/

OpenStudy (aravindg):

exactly :) proceed

OpenStudy (anonymous):

;O that was right?!

OpenStudy (aravindg):

cntinue :)

OpenStudy (anonymous):

V'=3a^2(d/dr)...?

OpenStudy (abb0t):

If your final is tmrw, I think u should focus on optimization and how to sketch a graph. Guarantee'd those will be worth more points.

OpenStudy (anonymous):

I'm going through everything on the study guide, I have until 7pm tomorrow.. it's 2am. thanks though. But, I have time to cover everything

OpenStudy (anonymous):

There will be no sketching a graph either. It's on a scantron.

OpenStudy (anonymous):

@AravindG did i do that wrong or?

OpenStudy (aravindg):

srry i was away ya u r right

OpenStudy (anonymous):

So now, I plug in right?

OpenStudy (aravindg):

yep

OpenStudy (anonymous):

a=30? d/dr=10?

OpenStudy (aravindg):

do u have final answr?

OpenStudy (anonymous):

I'm not getting the right answer

OpenStudy (aravindg):

whats the answr u have?

OpenStudy (anonymous):

3(30^2)(10)=27,000

OpenStudy (aravindg):

hey u dont have dr/dt !!

OpenStudy (anonymous):

V'=10?

OpenStudy (aravindg):

3a^2(d/dr(t))=10 from that find dr/dt

OpenStudy (anonymous):

1/270?

OpenStudy (aravindg):

now using that da/dt and the diff of area ie 2adr/dt

OpenStudy (aravindg):

2ada/dt will give u the answer :) srry gtg

OpenStudy (anonymous):

.... I'm so lost

OpenStudy (aravindg):

i think @Outkast3r09 will take over bye :)

OpenStudy (anonymous):

alright first figure out what you're trying to find the rate of You're looking for the surface area and you're given the rate of volume... you have to write Surface area in terms of V

OpenStudy (anonymous):

so what is the equation for surface area of cubes?

OpenStudy (anonymous):

omg!..... 6a^2

OpenStudy (anonymous):

SA=6a^2

OpenStudy (anonymous):

alright so \[SA=6a^2\] however you ned this equation for volume hmmmmm

OpenStudy (anonymous):

volume is \[x^3\] \[v=x^3\] if we solve for one side you get